### v1

# Theory for the spectral splitting exponent of exceptional points

Abstract
Exceptional points (EPs), singularities in non-Hermitian systems where eigenvalues and eigenstates coalesce, exhibit a dramatically enhanced response to perturbations compared to Hermitian degeneracies. This makes them exceptional candidates for sensing applications. The spectral splitting of an $N$th-order EP scales with perturbation strength $\epsilon$ over a wide range, from $\epsilon$ to $\epsilon^{1/N}$. Although the exact scaling exponent can be determined in principle by solving the characteristic equation, this approach becomes analytically intractable for large $N$ and often fails to yield useful physical insight. In this work, we develop a theory to directly predict the scaling exponent from the matrix positions of the perturbation. By using the Jordan block structure of the unperturbed Hamiltonian, we show that the splitting exponent can be analytically determined when the matrix positions of the perturbation satisfy some specific conditions. Our analytical framework provides a useful design principle for engineering perturbations to achieve a desired spectral response, facilitating the development of EP-based sensors.
I Introduction
Recently, non-Hermitian Hamiltonians have garnered considerable attention [1, 2] as an effective framework for describing open quantum and classical systems. In such systems, the effective Hamiltonian becomes non-Hermitian due to gain and loss [3, 4]. Notably, the introduction of non-Hermiticity gives rise to numerous phenomena without Hermitian counterparts, such as the emergence of exceptional points (EPs) [5, 6], the non-Bloch band theory [7, 8, 9, 10, 11, 12, 13, 14], non-Hermitian skin effects [7, 8, 15, 16, 17, 18, 12, 19, 14, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], novel non-Hermitian topological phases [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 35, 48, 36, 37], edge bursts [49, 50, 51], non-Bloch $\mathcal{PT}$ symmetry breaking [52, 53], and non-Hermitian invisibility [54, 55].
EPs correspond to degeneracies in non-Hermitian energy spectra [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76]. These spectral singularities give rise to a plethora of exotic physical phenomena, such as unidirectional invisibility [77, 78], topological energy transfer [79, 80] and laser mode selectivity [81, 82]. Due to the rich physics they enable, EPs have been realized experimentally across numerous platforms, including photonic crystals [83, 84, 85, 86], exciton-polariton systems [87], cavities [88, 89], metamaterials [90] and synthetic systems [91].
A salient property of EPs is their sensitive spectral response to perturbations, which is rooted in their singular wave-function structure. Unlike Hermitian degeneracies, where eigenstates remain orthogonal, EPs are characterized by the simultaneous coalescence of both eigenvalues and their corresponding eigenstates. This fundamental distinction leads to significantly different responses to perturbations. Specifically, under a perturbation of strength $\epsilon$, the spectral splitting for a Hermitian degeneracy scales at most linearly with $\epsilon$ ($\Delta E\sim\epsilon$). In stark contrast, the splitting for an $N$th-order EP (where $N$ eigenstates coalesce) can scale as $\Delta E\sim\epsilon^{1/N}$ [92, 93, 94]. The ratio of these responses, $\epsilon^{1/N}/\epsilon=\epsilon^{(1-N)/N}$, diverges in the limit $\epsilon\rightarrow 0$. This indicates that the response of an EP to a perturbation can be much stronger than that of any Hermitian degeneracy, a property that makes EPs exceptional candidates for enhancing sensitivity in sensing applications [95, 96, 93, 94, 97, 98, 99, 100, 101, 102].
While the strongest response, $\Delta E\sim\epsilon^{1/N}$, should in principle yield optimal sensor performance, this scaling is not guaranteed for an arbitrary perturbation [103, 104, 105]. For an $N$th-order EP, the exponent $\alpha$ characterizing the leading-order splitting (defined by $\Delta E\sim\epsilon^{\alpha}$) can range from $\alpha=1$ to $\alpha=1/N$. Specially, the exponent takes the general form $\alpha=1/k$, where $k$ is an integer satisfying $1\leq k\leq N$. The integer $k$ corresponds to the number of eigenstates that escape from the coalescence. Although the exponent $\alpha$ can always be determined in principle by solving the characteristic equation, this approach becomes analytically intractable for large $N$. Furthermore, numerical solutions, while feasible, often fail to provide physical insight into the relationship between the perturbation’s form and the resulting exponent. Therefore, an analytical theory capable of predicting $\alpha$ without solving for the full spectrum would be highly valuable, as it would enable the design of perturbations that yield a desired splitting scaling.
In a previous work, we explored the spectral splitting of infernal points—a class of EP whose order scales with system size—under perturbations [69]. Because the Hamiltonian at these points takes a Jordan form, we were able to analytically derive the condition for achieving the strongest possible response ($\alpha=1/N$). Intriguingly, the condition exhibits a simple relationship to the perturbation term’s positions in the matrix representation. Given that the Jordan block structure is a universal feature of Hamiltonians at EPs, we here extend this theory to general EPs of arbitrary order. We demonstrate that the splitting exponent $\alpha$ for a generic response can be determined analytically when the perturbation matrix satisfies specific algebraic conditions. This establishes a direct connection between the perturbation’s matrix structure and the exponent $\alpha$, with our previous result emerging as a special case [69].
The rest of this paper is organized as follows. In Sec. II, we introduce our general theoretical framework. In Sec. III, we analyze four distinct classes of perturbations and derive their characteristic spectral splitting behavior. In Sec. IV, we demonstrate the validity of our general framework using a fourth-order EP as a concrete example. Finally, we present a discussion and our conclusions in Sec. V.
II General framework
We consider a general $N\times N$ non-Hermitian Hamiltonian $H_{\rm N}$ which supports an $N$th-order EP for appropriate parameters. Without loss of generality, we assume the EP occurs at energy $E_{0}$. At this degeneracy, $H_{\rm N}$ possesses a single eigenstate, $|u_{0}\rangle$, satisfying $H_{\rm N}|u_{0}\rangle=E_{0}|u_{0}\rangle$. The complete $N$-dimensional Hilbert space is spanned by the Jordan chain, comprising this eigenstate and its N1N-1 associated vectors: $|u_{0}\rangle,|u_{1}\rangle,\cdots,|u_{N-1}\rangle$}. Correspondingly, the system has a single left eigenstate, $|v_{0}\rangle$, satisfying $H_{\rm N}^{\dagger}|v_{0}\rangle=E_{0}^{*}|v_{0}\rangle$, along with N1N-1 left associated vectors, $|v_{1}\rangle,\cdots,|v_{N-1}\rangle$. Owing to the Hamiltonian’s defectiveness, the standard biorthogonality condition $\langle v_{i}|u_{j}\rangle=\delta_{ij}$ breaks down. Instead, the vectors obey the Jordan chain structure [106], characterized by the relation $\langle v_{i}|u_{j}\rangle=\delta_{i+j,N-1}$ (A brief review is provided in Appendix A).
We introduce a perturbation to the system, $H_{p}=\epsilon H_{1}$, where $\epsilon$ is a dimensionless parameter that characterizes the perturbation strength. Under this perturbation, we assume that the original eigenenergy $E_{0}$ shifts to $E_{0}+\lambda$, and the modified eigenstate $|u\rangle$ can be expanded as
$|u\rangle=|u_{0}\rangle+\sum_{i=1}^{N-1}s_{i}|u_{i}\rangle,$ (1)
where $s_{i}$ denotes coefficients. Accordingly, the Schrödinger equation for the perturbed system is
$\left[H_{\rm N}+H_{p}-E_{0}-\lambda\right]|u\rangle=0.$ (2)
Substituting Eq. (1) into Eq. (2) and applying the Jordan chain structure from Eq. (A1), we obtain
$\sum_{i=1}^{N-1}s_{i}|u_{i-1}\rangle+\left(\epsilon H_{1}-\lambda\right)\left[|u_{0}\rangle+\sum_{i=1}^{N-1}s_{i}|u_{i}\rangle\right]=0.$ (3)
Left-multiplying Eq. (3) successively by $\langle v_{0}|$, $\langle v_{1}|$, $\cdots$, $\langle v_{L-1}|$ and applying the orthogonal normalization condition given in Eq. (28), we obtain
$\begin{split}\epsilon\left[H_{1}^{0,0}+\sum_{i=1}^{N-1}s_{i}H_{1}^{0,i}\right]-\lambda s_{N-1}&=0,\\ s_{N-1}+\epsilon\left[H_{1}^{1,0}+\sum_{i=1}^{N-1}s_{i}H_{1}^{1,i}\right]-\lambda s_{N-2}&=0,\\ s_{N-2}+\epsilon\left[H_{1}^{2,0}+\sum_{i=1}^{N-1}s_{i}H_{1}^{2,i}\right]-\lambda s_{N-3}&=0,\\ \vdots\qquad\qquad\qquad\ \quad&\\ s_{2}+\epsilon\left[H_{1}^{N-2,0}+\sum_{i=1}^{N-1}s_{i}H_{1}^{N-2,i}\right]-\lambda s_{1}&=0,\\ s_{1}+\epsilon\left[H_{1}^{N-1,0}+\sum_{i=1}^{N-1}s_{i}H_{1}^{N-1,i}\right]-\lambda\ &=0,\end{split}$ (4)
where $H_{1}^{i,j}$ represent the element $\langle v_{i}|H_{1}|u_{j}\rangle$. These equations can be written in compact form as:
$\left(\mathcal{J}_{\rm N}+\epsilon\mathcal{H}_{1}\right)S=\lambda S,$ (5)
where
$\mathcal{J}_{\rm N}=\begin{pmatrix}0&1&&&\\ &0&1&&\\ &&\ddots&\ddots&\\ &&&0&1\\ &&&&0\end{pmatrix}_{N\times N}$ (6)
is the $N\times N$ Jordan block,
$\mathcal{H}_{1}=\begin{pmatrix}H_{1}^{N-1,0}&H_{1}^{N-1,1}&\cdots&H_{1}^{N-1,N-1}\\ H_{1}^{N-2,0}&H_{1}^{N-2,1}&\cdots&H_{1}^{N-2,N-1}\\ \vdots&\vdots&\ddots&\vdots\\ H_{1}^{0,0}&H_{1}^{0,1}&\cdots&H_{1}^{0,N-1}\end{pmatrix}_{N\times N}$ (7)
and $S=\left(1,s_{1},s_{2},\cdots,s_{N-1}\right)^{T}$. Hence, solving the characteristic equation given in Eq. (2) is equivalent to solving the eigenvalues and eigenvectors of matrix $\mathcal{J}+\epsilon\mathcal{H}_{1}$.
III Analytical determination of the splitting exponent
Although our approach may appear to merely recast the characteristic equation into a different form, we will demonstrate its utility in determining the splitting exponent. This is possible when the perturbation matrix structure satisfies specific conditions, leveraging the unique properties of the Jordan block structure.
III.1 Case 1
The first case we consider corresponds to the situation where all matrix elements $H_{1}^{p,q}$ vanish for $p+q<N$ in $\mathcal{H}_{1}$. For this case, the general form of the perturbation matrix $\mathcal{H}_{1}$ reduces to
$\mathcal{H}_{1}=\begin{pmatrix}0&H_{1}^{N-1,1}&H_{1}^{N-1,2}&\cdots&H_{1}^{N-1,N-1}\\ 0&0&H_{2}^{N-1,2}&\cdots&H_{1}^{N-2,N-1}\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&0&\cdots&\ddots&H_{1}^{1,N-1}\\ 0&0&\cdots&\cdots&0\end{pmatrix}_{N\times N}.$ (8)
The matrix $\mathcal{J}_{\rm N}+\epsilon\mathcal{H}_{1}-\lambda\mathbbm{1}$ is evidently an upper triangular matrix with all diagonal elements equal to $\lambda$. Consequently, its characteristic polynomial is simply
$p\left(\lambda,\epsilon,N\right)=\det\left[\mathcal{J}_{\rm N}+\epsilon\mathcal{H}_{1}-\lambda\mathbbm{1}\right]=\left(-\lambda\right)^{N}.$ (9)
The unique root of $p(\lambda,\epsilon,N)$ is $\lambda=0$, indicating that the perturbation does not shift or split the EP. While this result may seem trivial, it leads to an important conclusion: not every perturbation can influence an EP.
III.2 Case 2
The second case we consider corresponds to the situation where only a single matrix element $H_{1}^{p,q}$ is non-vanishing. Based on the conclusion from Case 1, perturbations with $p+q\geqslant N$ leave the EP intact. For $p+q<N$, the eigenenergy shift $\lambda$ satisfies (demonstration is provided in Appendix B)
$\lambda^{N-p}-\epsilon H_{1}^{p,q}\lambda^{q}=0.$ (10)
The nonzero solutions are given by
$\lambda_{m}=\epsilon^{\frac{1}{N-p-q}}\left(H_{1}^{p,q}\right)^{\frac{1}{N-p-q}}e^{\frac{2im\pi}{N-p-q}},$ (11)
where $m=1,2,\cdots,N-p-q$. This result indicates that the spectral splitting scales as $\Delta E\sim\epsilon^{1/(N-p-q)}$. Given that $0\leqslant p+q<N$, a single-element perturbation can produce any splitting exponent of the form $\alpha=1/k$ for integers $k$ satisfying $1\leqslant k\leqslant N$. In particular, the choice $p=q=0$ yields the strongest possible scaling, $\Delta E\sim\epsilon^{1/N}$, which recovers the result from our previous work [69].
III.3 Case 3
The third case we consider corresponds to diagonal perturbations, meaning that the only nonzero elements of $\mathcal{H}_{1}$ are those whose indices satisfy p+q=N1p+q=N-1. Accordingly, the matrix $\mathcal{J}+\epsilon\mathcal{H}_{1}$ takes the form
$\mathcal{J}_{\rm N}+\epsilon\mathcal{H}_{1}=\begin{pmatrix}\epsilon H_{1}^{N-1,0}&\ 1\ &&&\\ &\epsilon H_{1}^{N-2,1}&\ 1\ &&\\ &&\ddots&\ \ddots\ &\\ &&&\ddots&\ 1\\ &&&&\epsilon H_{1}^{0,N-1}\end{pmatrix}_{N\times N}.$ (12)
Apparently, the characteristic polynomial of this case is simply
$p\left(\lambda,\epsilon,N\right)=\prod_{i=m}^{N}\left(\epsilon H_{1}^{N-m,m-1}-\lambda\right).$ (13)
The solutions are
$\lambda_{m}=\epsilon H_{1}^{N-m,m-1}\qquad(m=1,2,\cdots,N).$ (14)
This result shows that the perturbation induces a linear shift in the eigenvalues. The spectral consequence depends on the matrix elements: identical elements $H_{1}^{N-m,m-1}$ produce a uniform energy shift, preserving the EP, while non-identical elements lift the degeneracy, reducing the EP’s order or completely splitting the EP.
III.4 Case 4
The last case we consider corresponds to the situation where the only nonzero elements lie on the line $p+q=j$. For $j\geqslant N$, the EP remains intact, as established in Case 1. For j=N1j=N-1, the conclusion from Case 3 indicates a linear splitting exponent ($\alpha=1$).
For (N1)/2<j<N1(N-1)/2<j<N-1, we find that it is difficult to solve the eigenenergies of the perturbed Hamiltonian. However, the case $j\leqslant(N-1)/2$ is analytically solvable. Under this condition, the spectral modifications are given by (demonstration is provided in Appendix C)
$\lambda_{m}=\epsilon^{\frac{1}{N-j}}\left(\sum_{k=0}^{j}H_{1}^{j-k,k}\right)^{\frac{1}{N-j}}e^{\frac{2im\pi}{N-j}},$ (15)
where $m=1,2,\cdots,N-j$. This agrees with the results of Case 2. An important conclusion from Eq. (15) is that all matrix elements $H_{1}^{p,q}$ with $p+q=j\leqslant(N-1)/2$ contribute to the spectral splitting at the same order.
The consistent result $\alpha=1/(N-p-q)$ across the last three cases leads us to conjecture that the splitting exponent is a function only of the index sum $p+q$. In other words, $\alpha$ is governed exclusively by the locations of the nonzero elements of the perturbation matrix.
IV Application in a concrete model
While many models have been proposed to realize EPs, we select a non-Hermitian supersymmetric (SUSY) array model [92] to illustrate our theoretical framework. This model is ideally suited for this purpose, as it allows for the construction of EPs of any order. Specially, the Hamiltonian for an $N$-site non-Hermitian SUSY array is
$\begin{split}\hat{H}_{\rm L}=&\sum_{m=1}^{N}\left[\omega_{0}+i\gamma\left(N+1-2m\right)\right]c_{m}^{\dagger}c_{m}\\ &+\sum_{m=1}^{N-1}J\sqrt{m(N-m)}\left(c_{m}^{\dagger}c_{m+1}+{\rm h.c.}\right).\end{split}$ (16)
This Hamiltonian has a physical interpretation as describing a spin-(N1)/2(N-1)/2 particle under the influence of an effective magnetic field given by $\mathbf{B}=(J,0,i\gamma)$. The characteristic energy $\omega_{0}$ corresponds to the on-resonator frequency of each resonator or waveguide, depending on the concrete physical realization of this model.
The eigenenergies of this Hamiltonian is simply
$E_{L,n}=\omega_{0}+n\sqrt{J^{2}-\gamma^{2}},$ (17)
where $n=-(N-1),-(N-3),\cdots,(N-3),(N-1)$. When $J=\gamma$, all eigenenergies coalesce at $\omega_{0}$, forming an $N$th-order EP.
For concreteness, we focus on the case $N=4$, where the Hamiltonian hosts a fourth-order EP at $J=\gamma$. To apply our framework, we first express the Hamiltonian in matrix form. Using the basis $\psi=(c_{1},c_{2},c_{3},c_{4})^{T}$, the Hamiltonian at the EP is $\hat{H}_{4}=\psi^{{\dagger}}H_{4}\psi$, where
$H_{4}=\begin{pmatrix}\omega_{0}+3iJ&\sqrt{3}J&0&0\\ \sqrt{3}J&\omega_{0}+iJ&2J&0\\ 0&2J&\omega_{0}-iJ&\sqrt{3}J\\ 0&0&\sqrt{3}J&\omega_{0}-3iJ\end{pmatrix}$. (18)
The Hamiltonian can be transformed into its Jordan form via a similarity transformation:
$\mathcal{J}_{4}=\begin{pmatrix}\omega_{0}&1&0&0\\ 0&\omega_{0}&1&0\\ 0&0&\omega_{0}&1\\ 0&0&0&\omega_{0}\end{pmatrix}$, (19)
where $\mathcal{J}_{4}=U_{4}^{-1}H_{4}U_{4}$. The transformation matrix $U_{4}$ and its inverse $U_{4}^{-1}$ are of the form:
$U_{4}=\begin{pmatrix}-i&-\frac{1}{J}&\frac{i}{2J^{2}}&\frac{1}{6J^{3}}\\ -\sqrt{3}&\frac{2i}{\sqrt{3}J}&\frac{1}{2\sqrt{3}J^{2}}&0\\ i\sqrt{3}&\frac{1}{\sqrt{3}J}&0&0\\ 1&0&0&0\end{pmatrix}$, (20)
$U_{4}^{-1}=\begin{pmatrix}0&0&0&1\\ 0&0&\sqrt{3}J&-3iJ\\ 0&2\sqrt{3}J^{2}&-4i\sqrt{3}J^{2}&-6J^{2}\\ 6J^{3}&-6i\sqrt{3}J^{3}&-6\sqrt{3}J^{3}&6iJ^{3}\end{pmatrix}$. (21)
These two matrices can be expressed as
$\begin{split}U_{4}&=\left(|u_{0}\rangle,|u_{1}\rangle,|u_{2}\rangle,|u_{3}\rangle\right),\\ U_{4}^{-1}&=\left(|v_{3}\rangle,|v_{2}\rangle,|v_{1}\rangle,|v_{0}\rangle\right)^{\dagger},\end{split}$ (22)
where $|u_{i}\rangle$ and $|v_{i}\rangle$ are the right and left eigenvectors (including associated vectors) of $H_{4}$, which together satisfy the orthogonal normalization condition in Eq. (28).
We now consider five distinct perturbations whose expressions display different levels of complexity. Their explicit expressions are given by
$\begin{split}\hat{H}_{p1}&=\epsilon\frac{J}{6}c_{1}^{\dagger}c_{4},\\ \hat{H}_{p2}&=\epsilon\frac{J}{4\sqrt{3}}\left(c_{1}^{\dagger}c_{3}+c_{2}^{\dagger}c_{4}\right),\\ \hat{H}_{p3}&=\epsilon J\left[\frac{3}{2}c_{1}^{\dagger}c_{4}+\frac{1}{2}c_{2}^{\dagger}c_{3}+\frac{i\sqrt{3}}{2}\left(c_{1}^{\dagger}c_{3}-c_{2}^{\dagger}c_{4}\right)\right],\\ \hat{H}_{p4}&=\epsilon J\left[c_{2}^{\dagger}c_{2}+c_{3}^{\dagger}c_{3}+\sqrt{3}\left(c_{1}^{\dagger}c_{3}+c_{2}^{\dagger}c_{4}+ic_{1}^{\dagger}c_{2}-ic_{3}^{\dagger}c_{4}\right)\right],\\ \hat{H}_{p5}&=\epsilon J^{2}\left[\left(ic_{3}^{\dagger}c_{4}-ic_{1}^{\dagger}c_{2}+\mathrm{h.c.}\right)-\left(c_{1}^{\dagger}c_{3}+c_{2}^{\dagger}c_{4}-\mathrm{h.c.}\right)\right].\end{split}$ (23)
Although the physical meaning of these perturbations is clear in this form, their associated spectral splitting exponent is not directly evident. To apply our theoretical framework, we transform their matrix representations into the basis of the Jordan form using the same similarity transformation. Specifically, we express the perturbation in the original basis as $\hat{H}_{pi}=\psi^{{\dagger}}H_{pi}\psi$ and then apply the transformation to obtain its representation in the Jordan basis: $\epsilon\mathcal{H}_{1}^{pi}=U_{4}^{-1}H_{pi}U_{4}$. The results are
[…]
According to the framework established in Sec. III, we classify each perturbation and determine its spectral splitting exponent:
(1) $\epsilon\mathcal{H}_{1}^{p1}$ (Case 2): Its sole nonzero matrix element satisfies $p+q=0$, directly yielding an exponent of $\alpha=1/4$.
(2) $\epsilon\mathcal{H}_{1}^{p2}$ (Case 4): The two nonzero elements lie on the line $p+q=1$, which immediately gives $\alpha=1/3$.
(3) $\epsilon\mathcal{H}_{1}^{p3}$ (Case 2): With its only nonzero element at $p+q=2$, the exponent is $\alpha=1/2$.
(4) $\epsilon\mathcal{H}_{1}^{p4}$ (Case 3): This diagonal perturbation produces a linear splitting, corresponding to $\alpha=1$.
(5) $\epsilon\mathcal{H}_{1}^{p5}$ (Case 1): Despite having many terms in the original basis, all its nonzero elements are in the upper triangle ($p+q\geq N$), resulting in no splitting of the EP ($\alpha$ is undefined).
To demonstrate the predictive power of our theoretical framework, we now calculate the explicit magnitude of the spectral splitting, not just the exponent, using the formulas derived in Sec. III. For the five perturbations, the induced spectral splittings are respectively
$\begin{split}|\Delta E|^{(1)}&=\left|\left(\epsilon J^{4}\right)^{\frac{1}{4}}\right|=\epsilon^{\frac{1}{4}}\left|J\right|,\\ |\Delta E|^{(2)}&=\left|\left[\epsilon\left(\frac{1}{2}J^{3}+\frac{1}{2}J^{3}\right)\right]^{\frac{1}{3}}\right|=\epsilon^{\frac{1}{3}}\left|J\right|,\\ |\Delta E|^{(3)}&=\left|\left(\epsilon J^{2}\right)^{\frac{1}{2}}\right|=\epsilon^{\frac{1}{2}}\left|J\right|,\\ |\Delta E|^{(4)}&=\epsilon\left|J\right|,\\ |\Delta E|^{(5)}&=0.\end{split}$ (25)
Setting $J=1$ for simplicity yields the successive results $\epsilon^{1/4}$, $\epsilon^{1/3}$, $\epsilon^{1/2}$, $\epsilon$, and 0. Fig. 1 compares these analytical predictions (solid curves) with exact numerical results (dots), showing excellent agreement and confirming the accuracy of our theory.
Figure 1: Absolute eigenenergy splitting $|\Delta E|$ for the 4-site non-Hermitian SUSY array under five different perturbations $H_{p1}$ to $H_{p5}$. Dots represent numerical results, while solid curves show the analytical predictions $|\Delta E|=\epsilon^{1/4}$, $\epsilon^{1/3}$, $\epsilon^{1/2}$, $\epsilon$, and 0, respectively. Parameters are $\omega_{0}=0$, $\gamma=1$, and $J=1$.
V Discussions and conclusions
EPs are fundamental spectral singularities in non-Hermitian systems, and understanding their response to perturbations is crucial for developing EP-based applications. At an EP, the Hamiltonian becomes defective, preventing full diagonalization and necessitating a description via its Jordan block structure. In this work, we have developed a general theoretical framework that leverages this Jordan form to analytically predict the spectral splitting exponent $\alpha$ for a perturbed EP. A key finding of our theory is that the exponent $\alpha$ appears to be governed solely by the positions of the nonzero elements in the perturbation matrix. We validated this framework through both analytical and numerical studies of a fourth-order EP under distinct perturbations. The excellent agreement between our analytical predictions and exact numerical results confirms the power and accuracy of our approach. Consequently, the theoretical framework developed here provides a powerful design principle for engineering perturbations to achieve a desired spectral response, promising wide applications in non-Hermitian physics.
VI Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 12174455) and Guangdong Basic and Applied Basic Research Foundation (Grant No.2023B1515040023)
Appendix A A brief review of the Jordan chain structure
Consider an $N$th-order EP at energy $E_{0}$, described by an $N\times N$ non-Hermitian Hamiltonian $H_{N}$. This Hamiltonian possesses a unique eigenstate $|u_{0}\rangle$ and, as shown in Ref. [106], N1N-1 associated vectors $\{|u_{1}\rangle,\dots,|u_{N-1}\rangle\}$ that together form a Jordan chain satisfying:
$\begin{split}\left(H_{\rm N}-E_{0}\right)&|u_{0}\rangle=0,\\ \left(H_{\rm N}-E_{0}\right)&|u_{1}\rangle=|u_{0}\rangle,\\ &\vdots\\ \left(H_{\rm N}-E_{0}\right)&|u_{N-1}\rangle=|u_{N-2}\rangle.\end{split}$ (26)
Similarly, the left eigenstate $|v_{0}\rangle$ and its N1N-1 left associated vectors $\{|v_{1}\rangle,|v_{2}\rangle,\cdots,|v_{N-1}\rangle\}$ together form a Jordan chain satisfying:
$\begin{split}\langle v_{0}|&\left(H_{\rm N}-E_{0}\right)=0,\\ \langle v_{1}|&\left(H_{\rm N}-E_{0}\right)=\langle v_{0}|,\\ &\quad\vdots\\ \langle v_{N-1}|&\left(H_{\rm N}-E_{0}\right)=\langle v_{N-2}|.\end{split}$ (27)
These two sets of states satisfy the orthogonal normalization condition:
$\langle v_{i}|u_{j}\rangle=\delta_{i+j,N-1},$ (28)
where $i$ and $j$ take values in $\{0,1,\cdots,N-1\}$.
Appendix B Derivation of Eq. (10)
When the matrix $\mathcal{H}_{1}$ has only one nonzero element $H_{1}^{p,q}$, the system of equations in Eq. (4) simplifies to
$\begin{split}-\lambda s_{N-1}&=0,\\ s_{N-1}-\lambda s_{N-2}&=0,\\ s_{N-2}-\lambda s_{N-3}&=0,\\ \vdots\qquad\quad&\\ s_{N-p+1}-\lambda s_{N-p}&=0,\\ s_{N-p}+\epsilon s_{q}H_{1}^{p,q}-\lambda s_{N-p-1}&=0,\\ s_{N-p-1}-\lambda s_{N-p-2}&=0\\ \vdots\qquad\quad&\\ s_{2}-\lambda s_{1}&=0,\\ s_{1}-\lambda&=0.\end{split}$ (29)
Here, we have set $s_{0}=1$. Obviously, the solutions for $s_{i}$ are given by
$s_{i}=\begin{cases}\lambda^{i}&\mathrm{for}\ 0\leqslant i\leqslant N-p-1\\ 0&\mathrm{for}\ N-p\leqslant i\leqslant N-1\end{cases}.$ (30)
If $p+q<N$ (i.e., q<Npq<N-p), then $s_{q}=\lambda^{q}$. Substituting this into the equation $s_{N-p}+\epsilon s_{q}H_{1}^{p,q}-\lambda s_{N-p-1}=0$, one immediately obtains
$\epsilon\lambda^{q}H_{1}^{p,q}-\lambda^{N-p}=0,$ (31)
which is identical to Eq. (10) in the main text.
Appendix C Derivation of Eq. (15)
In Case 4, when j<N1j<N-1, Eq. (4) can be reduced as
$\begin{split}\epsilon s_{j}H_{1}^{0,j}-\lambda s_{N-1}&=0,\\ s_{N-1}+\epsilon s_{j-1}H_{1}^{1,j-1}-\lambda s_{N-2}&=0\\ \vdots\qquad&\\ s_{N-j+1}+\epsilon s_{1}H_{1}^{j-1,1}-\lambda s_{N-j}&=0\\ s_{N-j}+\epsilon H_{1}^{j,0}-\lambda s_{N-j-1}&=0,\\ s_{N-j-1}-\lambda s_{N-j-2}&=0,\\ \vdots\qquad&\\ s_{2}-\lambda s_{1}&=0,\\ s_{1}-\lambda&=0.\end{split}$ (32)
Similarly, it is easy to find $s_{i}=\lambda^{i}$ for $i=\{0,1,\cdots,N-j-1\}$. If a more stringent constraint, $j\leqslant\frac{1}{2}\left(N-1\right)$, is put, then $s_{i}$ for $i\geqslant N-j$ can be represented as
$\begin{split}s_{N-j}=&\lambda^{N-j}+\epsilon H_{1}^{j,0},\\ s_{N-j+1}=&\lambda\left(\lambda^{N-j}+\epsilon H_{1}^{j,0}+\epsilon H_{1}^{j-1,1}\right),\\ s_{N-j+2}=&\lambda^{2}\left(\lambda^{N-j}+\epsilon H_{1}^{j,0}+\epsilon H_{1}^{j-1,1}+\epsilon H_{1}^{j-2,2}\right),\\ \vdots\\ s_{N-1}=&\lambda^{j-1}\left(\lambda^{N-j}-\epsilon\sum_{k=0}^{j-1}H_{1}^{j-k,k}\right).\end{split}$ (33)
Substituting the last equation of (33) into the first equation of (32) yields
$\lambda^{j}\left(\lambda^{N-j}-\epsilon\sum_{k=0}^{j}H_{1}^{j-k,k}\right)=0.$ (34)
The nonzero solutions of this equation are
$\lambda_{m}=\epsilon^{\frac{1}{N-j}}\left(\sum_{k=0}^{j}H_{1}^{j-k,k}\right)^{\frac{1}{N-j}}e^{\frac{2im\pi}{N-j}},$ (35)
for $m=1,2,\cdots,N-j$, which is the result presented in Eq. (15).

### v2

@@
 I Introduction
-Recently, non-Hermitian Hamiltonians have garnered considerable attention [1, 2] as an effective framework for describing open quantum and classical systems.
-In such systems, the effective Hamiltonian becomes non-Hermitian due to gain and loss [3, 4].
-Notably, the introduction of non-Hermiticity gives rise to numerous phenomena without Hermitian counterparts, such as the emergence of exceptional points (EPs) [5, 6], the non-Bloch band theory [7, 8, 9, 10, 11, 12, 13, 14], non-Hermitian skin effects [7, 8, 15, 16, 17, 18, 12, 19, 14, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], novel non-Hermitian topological phases [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 35, 48, 36, 37], edge bursts [49, 50, 51], non-Bloch $\mathcal{PT}$ symmetry breaking [52, 53], and non-Hermitian invisibility [54, 55].
-EPs correspond to degeneracies in non-Hermitian energy spectra [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76].
-These spectral singularities give rise to a plethora of exotic physical phenomena, such as unidirectional invisibility [77, 78], topological energy transfer [79, 80] and laser mode selectivity [81, 82].
-Due to the rich physics they enable, EPs have been realized experimentally across numerous platforms, including photonic crystals [83, 84, 85, 86], exciton-polariton systems [87], cavities [88, 89], metamaterials [90] and synthetic systems [91].
+Recently, non-Hermitian Hamiltonians have garnered considerable attention [RevModPhys.93.015005, ashida2020non] as an effective framework for describing open quantum and classical systems.
+In such systems, the effective Hamiltonian becomes non-Hermitian due to gain and loss [Rotter_2009, PhysRevLett.70.2273].
+Notably, the introduction of non-Hermiticity gives rise to numerous phenomena without Hermitian counterparts, such as the emergence of exceptional points (EPs) [W.D.Heiss_2004, berry2004physics], the non-Bloch band theory [PhysRevLett.121.086803, PhysRevLett.123.066404, PhysRevB.101.195147, PhysRevX.14.021011, HU202551, zhang2022universal, PhysRevLett.133.216401, PhysRevB.107.115412], non-Hermitian skin effects [PhysRevLett.121.086803, PhysRevLett.123.066404, PhysRevLett.124.056802, PhysRevLett.125.126402, PhysRevLett.124.086801, PhysRevLett.125.226402, zhang2022universal, PhysRevLett.125.186802, PhysRevB.107.115412, PhysRevB.109.L081108, PhysRevB.108.L060204, PhysRevB.106.L241112, li2021impurity, PhysRevResearch.5.033058, PhysRevB.108.075122, PhysRevB.108.205423, PhysRevLett.127.256402, PhysRevLett.131.116601, PhysRevLett.127.116801, PhysRevB.111.L140201, li2020critical, PhysRevB.104.165117, PhysRevB.107.155430, li2024observation, PhysRevLett.123.016805, PhysRevB.102.241202, PhysRevB.103.045420, Chen_2024, manna2023inner], novel non-Hermitian topological phases [PhysRevX.8.031079, PhysRevX.9.041015, PhysRevLett.120.146402, PhysRevB.110.L121401, PhysRevB.99.201103, PhysRevLett.116.133903, PhysRevLett.121.136802, PhysRevLett.126.216405, PhysRevLett.122.076801, PhysRevB.102.205118, PhysRevLett.123.016805, PhysRevB.99.081302, PhysRevB.102.241202, PhysRevB.103.045420, PhysRevB.111.085148, PhysRevB.111.094109, PhysRevB.111.195419], edge bursts [PhysRevLett.128.120401, PhysRevB.108.235422, PhysRevLett.133.070801], non-Bloch $\mathcal{PT}$ symmetry breaking [PhysRevLett.132.050402, PhysRevLett.126.230402], and non-Hermitian invisibility [PhysRevA.82.032111, PhysRevB.106.094205].
+EPs correspond to degeneracies in non-Hermitian energy spectra [Heiss_2012, ding2022non, PhysRevResearch.2.043045, PhysRevLett.126.086401, PhysRevB.105.075420, PhysRevB.104.L201104, PhysRevB.104.L161116, PhysRevB.104.L161117, PhysRevLett.123.066405, gohsrich2024, PhysRevLett.128.010402, denner2021exceptional, PhysRevB.110.L201104, wang2025, PhysRevResearch.6.043023, PhysRevLett.127.186602, PhysRevLett.118.045701, PhysRevB.99.081102, PhysRevA.98.042114, PhysRevB.99.161115, PhysRevLett.126.010401, PhysRevB.110.L020104, PhysRevLett.134.133801, 18gg-gvzc, PhysRevResearch.4.023130, bid2025fragmentedexceptionalpointsbulk, PhysRevResearch.7.023062], and share intriguing connections with Newton polygons and tropical geometry [jaiswal2023characterizing, banerjee2023tropical].
+These spectral singularities give rise to a plethora of exotic physical phenomena, such as unidirectional invisibility [Lin2011unidirectional, Peng2014], topological energy transfer [Doppler2016, Xu2016] and laser mode selectivity [Hossein2014, Feng2014].
+Due to the rich physics they enable, EPs have been realized experimentally across numerous platforms, including photonic crystals [ozdemir2019parity, regensburger2012parity, zhou2023observation, kim2016direct], exciton-polariton systems [gao2015observation], cavities [PhysRevX.6.021007, PhysRevLett.103.134101], metamaterials [Park2020] and synthetic systems [yang2023realization].
 A salient property of EPs is their sensitive spectral response to perturbations, which is rooted in their singular wave-function structure.
@@
 Specifically, under a perturbation of strength $\epsilon$, the spectral splitting for a Hermitian degeneracy scales at most linearly with $\epsilon$ ($\Delta E\sim\epsilon$).
-In stark contrast, the splitting for an $N$th-order EP (where $N$ eigenstates coalesce) can scale as $\Delta E\sim\epsilon^{1/N}$ [92, 93, 94].
+In stark contrast, the splitting for an $N$th-order EP (where $N$ eigenstates coalesce) can scale as $\Delta E\sim\epsilon^{1/N}$ [PhysRevA.101.033820, chen2017exceptional, hodaei2017enhanced].
 The ratio of these responses, $\epsilon^{1/N}/\epsilon=\epsilon^{(1-N)/N}$, diverges in the limit $\epsilon\rightarrow 0$.
-This indicates that the response of an EP to a perturbation can be much stronger than that of any Hermitian degeneracy, a property that makes EPs exceptional candidates for enhancing sensitivity in sensing applications [95, 96, 93, 94, 97, 98, 99, 100, 101, 102].
-While the strongest response, $\Delta E\sim\epsilon^{1/N}$, should in principle yield optimal sensor performance, this scaling is not guaranteed for an arbitrary perturbation [103, 104, 105].
+This indicates that the response of an EP to a perturbation can be much stronger than that of any Hermitian degeneracy, a property that makes EPs exceptional candidates for enhancing sensitivity in sensing applications [PhysRevLett.112.203901, PhysRevA.93.033809, chen2017exceptional, hodaei2017enhanced, PhysRevLett.123.213901, PhysRevLett.122.153902, PhysRevApplied.12.024002, PhysRevLett.125.240506, PhysRevA.101.053846, PhysRevResearch.6.033284].
+The ultimate performance in practical devices, however, is limited by noise, which bounds the signal-to-noise ratio [lau2018fundamental, Chen_2019].
+Strategies to overcome this limitation remain an active area of research [Kononchuk2022].
+A deeper understanding of the spectral splitting behavior can shed new light on these key problems.
+While the strongest response, $\Delta E\sim\epsilon^{1/N}$, should in principle yield optimal sensor performance, this scaling is not guaranteed for an arbitrary perturbation [PhysRevLett.127.186601, PhysRevLett.134.153601, wu2021high, PhysRevB.99.241403].
 For an $N$th-order EP, the exponent $\alpha$ characterizing the leading-order splitting (defined by $\Delta E\sim\epsilon^{\alpha}$) can range from $\alpha=1$ to $\alpha=1/N$.
@@
 Therefore, an analytical theory capable of predicting $\alpha$ without solving for the full spectrum would be highly valuable, as it would enable the design of perturbations that yield a desired splitting scaling.
-In a previous work, we explored the spectral splitting of infernal points—a class of EP whose order scales with system size—under perturbations [69].
+In a previous work, we explored the spectral splitting of infernal points—a class of EP whose order scales with system size—under perturbations [wang2025].
 Because the Hamiltonian at these points takes a Jordan form, we were able to analytically derive the condition for achieving the strongest possible response ($\alpha=1/N$).
@@
 We demonstrate that the splitting exponent $\alpha$ for a generic response can be determined analytically when the perturbation matrix satisfies specific algebraic conditions.
-This establishes a direct connection between the perturbation’s matrix structure and the exponent $\alpha$, with our previous result emerging as a special case [69].
+This establishes a direct connection between the perturbation’s matrix structure and the exponent $\alpha$, with our previous result emerging as a special case [wang2025].
 The rest of this paper is organized as follows.
@@
 Owing to the Hamiltonian’s defectiveness, the standard biorthogonality condition $\langle v_{i}|u_{j}\rangle=\delta_{ij}$ breaks down.
-Instead, the vectors obey the Jordan chain structure [106], characterized by the relation $\langle v_{i}|u_{j}\rangle=\delta_{i+j,N-1}$ (A brief review is provided in Appendix A).
+Instead, the vectors obey the Jordan chain structure [seyranian2003multiparameter], characterized by the relation $\langle v_{i}|u_{j}\rangle=\delta_{i+j,N-1}$ (A brief review is provided in Appendix A).
 We introduce a perturbation to the system, $H_{p}=\epsilon H_{1}$, where $\epsilon$ is a dimensionless parameter that characterizes the perturbation strength.
@@
 Given that $0\leqslant p+q<N$, a single-element perturbation can produce any splitting exponent of the form $\alpha=1/k$ for integers $k$ satisfying $1\leqslant k\leqslant N$.
-In particular, the choice $p=q=0$ yields the strongest possible scaling, $\Delta E\sim\epsilon^{1/N}$, which recovers the result from our previous work [69].
+In particular, the choice $p=q=0$ yields the strongest possible scaling, $\Delta E\sim\epsilon^{1/N}$, which recovers the result from our previous work [wang2025].
 III.3 Case 3
@@
 (15) is that all matrix elements $H_{1}^{p,q}$ with $p+q=j\leqslant(N-1)/2$ contribute to the spectral splitting at the same order.
-The consistent result $\alpha=1/(N-p-q)$ across the last three cases leads us to conjecture that the splitting exponent is a function only of the index sum $p+q$.
+The consistent result $\alpha=1/(N-p-q)$ across the last three cases leads us to conjecture that the splitting exponent is a function only of the index sum $p+q$, i.e., $j$.
+When the perturbation matrix contains terms with different values of $j$, the minimum value of $j$ governs the splitting exponent in the extremely weak-perturbation regime.
 In other words, $\alpha$ is governed exclusively by the locations of the nonzero elements of the perturbation matrix.
+Numerical verification of this conjecture is provided in Appendix D for the model discussed in Section $\mathbf{IV}$.
 IV Application in a concrete model
-While many models have been proposed to realize EPs, we select a non-Hermitian supersymmetric (SUSY) array model [92] to illustrate our theoretical framework.
+While many models have been proposed to realize EPs, we select a non-Hermitian supersymmetric (SUSY) array model [PhysRevA.101.033820] to illustrate our theoretical framework.
 This model is ideally suited for this purpose, as it allows for the construction of EPs of any order.
@@
 Figure 1: Absolute eigenenergy splitting $|\Delta E|$ for the 4-site non-Hermitian SUSY array under five different perturbations $H_{p1}$ to $H_{p5}$.
-Dots represent numerical results, while solid curves show the analytical predictions $|\Delta E|=\epsilon^{1/4}$, $\epsilon^{1/3}$, $\epsilon^{1/2}$, $\epsilon$, and 0, respectively.
+The blue, magenta, green, orange and red dots represent numerical results for $H_{p1}$, $H_{p2}$, $H_{p3}$, $H_{p4}$ and $H_{p5}$, respectively.
+Solid curves crossing the dots correspond to the analytical predictions $|\Delta E|=\epsilon^{1/4}$, $\epsilon^{1/3}$, $\epsilon^{1/2}$, $\epsilon$, and 0, respectively.
 Parameters are $\omega_{0}=0$, $\gamma=1$, and $J=1$.
@@
 This Hamiltonian possesses a unique eigenstate $|u_{0}\rangle$ and, as shown in Ref.
-[106], N1N-1 associated vectors $\{|u_{1}\rangle,\dots,|u_{N-1}\rangle\}$ that together form a Jordan chain satisfying:
+[seyranian2003multiparameter], N1N-1 associated vectors $\{|u_{1}\rangle,\dots,|u_{N-1}\rangle\}$ that together form a Jordan chain satisfying:
 $\begin{split}\left(H_{\rm N}-E_{0}\right)&|u_{0}\rangle=0,\\ \left(H_{\rm N}-E_{0}\right)&|u_{1}\rangle=|u_{0}\rangle,\\ &\vdots\\ \left(H_{\rm N}-E_{0}\right)&|u_{N-1}\rangle=|u_{N-2}\rangle.\end{split}$ (26)
@@
 for $m=1,2,\cdots,N-j$, which is the result presented in Eq. (15).
+When $\frac{1}{2}(N-1)<j<N-1$, or equivalently Nj1<j<N1N-j-1<j<N-1, there exists an integer $t<j$ such that $s_{i}\not=\lambda^{i}$ for $i=j,j-1,\cdots,t$.
+This means that Eq.
+(32) cannot be reduced to Eq.
+(33).
+Consequently, it is difficult to solve Eq.
+(32) analytically.
+Numerical results in appendix D show that a perturbation in this regime typically induces multiple split eigenvalues, all of which scale as $\epsilon^{\frac{1}{N-j}}$.
+Appendix D Perturbations beyond the four types discussed in section $\mathbf{III}$
+In Section $\mathbf{III}$, we discuss four types of perturbations that can be solved analytically.
+For more general cases, we conjecture that the spectral splitting exponent is determined by the smallest index $j$ among the nonzero elements of the perturbation matrix, because if we consider all non-zero elements separately, those with the smallest $j$ obviously governs the spectral splitting.
+Figure 2: Spectral splitting induced by (a) $\hat{H}_{p6}$ and (b) $\hat{H}_{p7}$.
+The orange dots represent numerical results.
+The three lines in (a) correspond to $\ln|\Delta E|=\frac{1}{4}\ln\epsilon+C$ with $C=0.07$, $0.008$, and 0.09-0.09, respectively.
+The two lines in (b) correspond to $\ln|\Delta E|=\frac{1}{2}\ln\epsilon+\frac{1}{2}\ln\frac{3\pm\sqrt{5}}{2}$.
+Parameters are $\omega_{0}=0$, $\gamma=1$, and $J=1$.
+In this section, we numerically verify this conjecture using the non-Hermitian SUSY array model from Section $\mathbf{IV}$.
+First, we investigate the case that the Hamiltonian $\hat{H}_{4}$ is perturbed by $\hat{H}_{p6}=\hat{H}_{p1}+\hat{H}_{p2}$.
+According to Eq.
+(24), the minimum index $j$ for nonzero elements in the perturbation matrix is $j_{min}=0$ .
+The numerical results in Fig.
+2 show that the splitting of eigenenergy for all three branches scales as $\epsilon^{\frac{1}{N-j_{min}}}=\epsilon^{\frac{1}{4}}$ for small $\epsilon$, confirming the predicted exponent.
+Next we examine the perturbation
+$\hat{H}_{p7}$ $=$ $\frac{\epsilon J}{2\sqrt{3}}\left[2\left(c_{1}^{\dagger}c_{2}+c_{3}^{\dagger}c_{4}\right)-i\left(c_{1}^{\dagger}c_{3}-c_{2}^{\dagger}c_{4}\right)\right$. (36)
+$\left.-\sqrt{3}\left(c_{1}^{\dagger}c_{4}-c_{2}^{\dagger}c_{3}\right)\right],$
+with its corresponding perturbation matrix given by
+$\epsilon\mathcal{H}_{1}^{p7}={\color[rgb]{0,0,0}\epsilon}\begin{pmatrix}0&0&0&0\\ J^{2}&0&0&0\\ 0&J^{2}&0&0\\ 0&0&J^{2}&0\end{pmatrix}$. (37)
+[…]